OpenStudy (anonymous):

Suppose $2,400 is invested in an account at an annual interest rate of 6.1% compounded continuously. How long (to the nearest tenth of a year) will it take the investment to double in size?

OpenStudy (anonymous):

Here we have to use the compound interest formulas . do you know it?

OpenStudy (anonymous):

yes 2400(1+0.61/12)^n*t?

OpenStudy (anonymous):

so this is the formula for interest and principal + interest = amount We have to find the time in which the principal gets double i.e. amount is 2 x 2400 = 4800

OpenStudy (anonymous):

\[e^{.061t}=2\] solve for \(t\) in two steps

OpenStudy (anonymous):

take the log, get \[.061t=\ln(2)\] divide by \(.061\) to get \[t=\frac{\ln(2)}{.061}\]

OpenStudy (anonymous):

forget the 24000 it is a red herring, has nothing to do with the problem doubling time is doubling time, no matter how much money you start with

OpenStudy (anonymous):

also it is continuous compounding, so you use \[A=P_0e^{rt}\]

OpenStudy (anonymous):

its just going to continue

OpenStudy (anonymous):

A = P + I 4800 = 2400 + 2400(1+6.1/100)^t? 2400 = 2400(1+6.1/100)^t t = 12 after rounding

OpenStudy (anonymous):

yes it does not matter whatever be the principal @satellite73 is all correct but i showed u to make formulas clear